1\x+21, x<0 -Sketch the graph of this piece-wise defined function: S(x) = {3 05x<2 1(x+1), x22
Answer #1

solution: Given functiun: - ter) = TH+21 x Lo ocx<2 l (a + 1)2 x>, 2 +(x) = x + 2) aco f(re) ii) f(x) = 3 fcal = 3 of a <2 f(۱۱۱) fx) = ( +1) 2 * ) 2 ادا f(re) کر م دیا ه ہے اf(x) = (x+1)2 f(x) = 3 f(x) = (x+21 -5 -4 -3 -2 -1 0 1 2 3 4